QUIZ TIME REMAININGCALCEXCELOVERVIEW HELP SUBMI QUIZ 1 2 3 4 5 6 7 8 9 Earl was engaged by Farm Corp. to perform consulting services. Earts compensation for these services consisted of 1,000 shares of Farms $10 par value common stock, to be issued to Earl on completion of Earls services. On the execution date of Earls employment contract, Farms stock had a market value of $40 per share. Six months later, when Earls services were completed and the stock issued, the stocks market value was $50 per share. Farms management estimated that Earls services were worth $100,000 in cost savings to the company. As a result of this transaction, additional paid-in capital should increase by $40,000 $90,000 $100,000 $30,000 650PM 11/4/2018 ^僱 O e Type here to search
Answer #1

Option D is the answer

The total value of compensation is based on the date the contract is executed.

Therefore total compensation cost = (1000*40) = 40,000

Par value of common stock = (1000*10) = 10,000

Increase in additional paid in capital = 30,000 (40,000-10,000)